OpenStudy (anonymous):

a culture started with 6000 bacteria. After 3 hours, it grew to 7200 bacteria. predict how many bacteria will be present after 19 hours. Round your answer to the nearst whole number Note:when solving for k, round to four decimal places.

OpenStudy (anonymous):

\[P=Ae ^{kt}\]

OpenStudy (anonymous):

that the equation given

OpenStudy (anonymous):

\[\frac{7200}{6000}=1.2\] i.e. it increases by 20% in three hours you can model this as \[6000\times (1.2)^{\frac{t}{3}}\]

OpenStudy (anonymous):

i have no idea how you got that

OpenStudy (anonymous):

then replace \(t\) by \(19\) and compute \[6000\times (1.2)^{\frac{19}{3}}\] with a calculator

OpenStudy (anonymous):

lets go slow because i only used the numbers you wrote

OpenStudy (anonymous):

you start at 6000 right?

OpenStudy (anonymous):

so what is the fracton on 1.2

OpenStudy (anonymous):

then it grew to 7200 in 3 hours

OpenStudy (anonymous):

that is an increase of 20% as \(7200\div 6000=1.2\)

OpenStudy (anonymous):

model as \[\large P(t)=6000\times (1.2)^{\frac{t}{3}}\]

OpenStudy (anonymous):

still have no idea do i multiply 6000 and 1.2

OpenStudy (anonymous):

to increase a number by 20% you multiply it by \(1.2\) this increases by 20% every three hours

OpenStudy (anonymous):

in three hours it is \(6000\times 1.2=7200\) in six hours it is \(6000\times 1.2^2\) in nine hours it is \(6000\times 1.2^3\)

OpenStudy (anonymous):

in 19 hours it is \[\large 6000\times (1.2)^{\frac{19}{3}}\]